# Cost-aware Stopping for Bayesian Optimization

### v3

@@ 3.1 Theoretical Guarantees on Cost-Adjusted Simple Regret
Let $X$ be compact, and let $f:X\to\mathbb{R}$ be a random function with constant mean $\mu$. Consider a Bayesian optimization algorithm that begins at some initial point $x_{1}\in X$ with cost $C=c(x_{1})$, acquires additional points using either the PBGI or LogEIPC acquisition function, and uses the PBGI/LogEIPC stopping rule. Let $\tau=\min_{t\geq 1}\{\sup_{x\in X}\alpha_{t}^{\mathrm{LogEIPC}}(x)<0\}$ be the algorithm’s stopping time, and denote the posterior expected improvement function by $\alpha_{t}^{\mathrm{EI}}(x)=\mathrm{EI}_{f\mid x_{1:t},y_{1:t}}\left(x;y^{*}_{1:t}\right)$. Then, $\alpha_{t}^{\mathrm{EI}}(x_{t+1})\geq c(x_{t+1})$ for all $t<\tau$.
Proofs are in Appendix˜B. As a consequence of this claim, we show in Theorem˜2 that the PBGI/LogEIPC stopping rule, when paired with PBGI or LogEIPC, achieve strictly better cost-adjusted simple regret than the naive strategy of stopping immediately after the initial evaluation. While this may sound obvious, many acquisition function–stopping rule pairs fail to satisfy this no-worse-than-immediate property in practice (see Figures˜2, 3 and 4), largely because most stopping rules are designed with simple regret alone in mind and do not explicitly account for evaluation costs. To our knowledge, this is the first result establishing a formal guarantee on cost-adjusted simple regret for Bayesian optimization, generalizing beyond simpler settings such as Pandora’s Box, which assumes independent and discrete evaluations.
@@
Consider the setting, acquisition function, and stopping rule specified in Lemma˜1. Let $U=\mu-\mathbb{E}[\min_{x\in X}f(x)]<\infty$, then the algorithm’s expected cost-adjusted simple regret is bounded by
@@ 3.1.1 Implication in the Budget-constrained Setting
We first note that in the discrete Pandora’s Box setting, under an expected budget constraint $B$, minimizing $\alpha_{t}^{\mathrm{PBGI}}(x)$ is Bayesian-optimal: it achieves the lowest expected simple regret, among all algorithms satisfying $\mathbb{E}\left[\sum_{t=1}^{\tau}c(x_{t})\right]\leq B$, and the cost function used in defining $\alpha_{t}^{\mathrm{PBGI}}(x)$ is $\lambda c(x)$ for some cost-scaling factor $\lambda$ which depends on $B$ (Xie et al., 2024, Theorem 2).
@@
Consider the setting, acquisition function, stopping rule, and notation specified in Lemmas˜1 and 2, but with costs rescaled by a factor $\lambda>0$: both acquisition values and stopping conditions are computed using $\lambda c(\cdot)$. If the cost-scaling factor is set to $\lambda=\frac{U}{B-C}$, then the algorithm’s expected cumulative unscaled cost satisfies $\mathbb{E}\!\left[\sum_{t=1}^{\tau}c(x_{t})\right]\leq B$.
If this choice of $\lambda$ proves overly conservative—leading to underspending within the budget—it can be paired with the PBGI-D variant of Xie et al. (2024), which starts with $\lambda=\lambda_{0}$ and halves it each time stopping is triggered. Choosing $\lambda_{0}=U/(B-C)$ aligns the initial fixed-$\lambda$ phase with the budget in expectation, while ensuring that the adaptive decay is activated as designed. In Figure˜11 in Appendix˜D, we show PBGI-D with this choice of $\lambda_{0}$ is competitive.
@@ 3.2 Practical Implementation Considerations for PBGI/LogEIPC Stopping Rule
In practice, evaluation costs are often not known in advance. They can be modeled either (i) deterministically, using domain knowledge (e.g., proportional to model size in hyperparameter tuning), or (ii) stochastically, via a Gaussian process over log costs. In Section˜4, we present empirical benchmark results under both modeling approaches. For the stochastic approach, we follow Astudillo et al. (2021, Proposition 2) to model $\ln c(x)$ as a Gaussian process with posterior mean $\mu_{\ln c}$ and variance $\sigma^{2}_{\ln c}$ . To compute LogEIPC or PBGI, as well as their related stopping rules (including prior acquisition-based ones and ours), we replace $c(x)$ in Equations˜4, 5 and 9 with $\mathbb{E}[c(x)]=\exp(\mu_{\ln c}(x)+(\sigma_{\ln c}(x))^{2}/2)$. See Section˜D.3 for further discussion.
@@
Although our stopping rule has theoretical guarantees, in practice, it—like other adaptive stopping rules—can still suffer from spurious stops caused by two main sources which we now discuss. First, early in the optimization process, the Gaussian process model parameters often fluctuate significantly as new data points are collected, causing unstable stopping signals. To mitigate this, we enforce a stabilization period consisting of the first several evaluations, during which we do not allow any stopping rule to trigger. Second, imperfect optimization of the acquisition function—which is especially common in higher-dimensional search spaces—can lead to misleading stopping signals. To handle this, we use a debounce strategy, requiring the stopping rule to consistently indicate stopping over several consecutive iterations before stopping optimization.
@@ 4 Experiments
To evaluate our proposed PBGI/LogEIPC stopping rule, we design three complementary sets of experiments that progressively test its performance. First, we consider an idealized, low-dimensional Bayesian regret setting in which the Gaussian process model exactly matches the true objective. This controlled environment allows us to isolate the effect of different stopping rules without interference from modeling or optimization errors. Then, we move to a higher-dimensional Bayesian regret setting where each acquisition-function minimization must be approximated via a numerical optimizer. Finally, we test in practical scenarios with potential objective model mismatch, using the LCBench hyperparameter tuning benchmark and the NATS neural architecture size search benchmark. In each case, we compare pairs of acquisition functions and stopping rules, which we describe next.
@@
We consider four common acquisition functions that were discussed in Section˜2: log expected improvement per cost (LogEIPC), Pandora’s Box Gittins Index (PBGI), lower confidence bound (LCB), and Thompson sampling (TS)—chosen for their competitive performance, computational efficiency, and close connections to existing stopping rules.
@@
We compare the proposed PBGI/LogEIPC stopping rule against several stopping rules from prior work: UCB–LCB Makarova et al. (2022), LogEIPC-med Ishibashi et al. (2023), SRGap-med Ishibashi et al. (2023), and PRB Wilson (2024). We also include two simple heuristics used in practice: Convergence and GSS. UCB–LCB stops once the gap between upper and lower confidence bounds falls below a configurable threshold $\theta$. LogEIPC-med stops when the log expected improvement per cost drops beneath $\log(\eta)$ plus the median of its initial $I$ values. SRGap-med stops when the gap of the expected minimum simple regret falls below $\chi$ times the median of its initial $I$ values. PRB triggers once a probabilistic regret bound satisfies the regret tolerance $\epsilon$ and confidence $\delta$ parameters. Convergence stops as soon as the best observed value remains unchanged for $w$ iterations. GSS stops if the recent improvement is less than $\phi\times\mathrm{IQR}$ over the past $w$ trials where $\mathrm{IQR}$ denotes the inter-quartile range of current observations. Finally, we include two reference baselines: Immediate, which stops right after the initial evaluation (see Section˜3.1), and Hindsight, which, for each trial, selects the stopping time that yields the lowest cost-adjusted simple regret in hindsight, thus providing a lower bound on achievable performance.
@@ 4.1 Bayesian Regret
In the 1-dimensional experiments, we perform an exhaustive grid search over $[0,1]$ to isolate the effect of stopping behavior from numerical optimization. In the 8-dimensional setting, due to instability from higher dimensionality, we apply moving average with a window size of 20 when computing the PBGI/LogEIPC stopping condition. See Figure˜5 in Appendix˜C for an illustration. More details on our experiment setup, and computational considerations are provided in Appendices˜C and D.
Figure˜2 shows a comparison of cost-adjusted regret for acquisition function and stopping rule pairings under different cost-scaling factors $\lambda$ in the 1-dimensional setting. From the plot, our PBGI/LogEIPC stopping rule pairing with LogEIPC or PBGI acquisition function delivers competitive performance for each $\lambda$, and is particularly strong in handling high-cost scenarios—those with large $\lambda$.
In the 1-dimensional setting, the Bayesian optimization problem is relatively straightforward and all acquisition functions attain nearly identical hindsight optima, independent of cost scale or cost type. In the 8-dimensional experiments, however, clear gaps emerge: Figure˜3 compares cost-adjusted regret under uniform, linear, and periodic cost regimes. Under uniform and linear costs, LogEIPC, PBGI, and LCB perform similarly and substantially outperform TS, while in the periodic case, LogEIPC and PBGI outperform LCB and TS. In every case, combining our stopping rule with the LogEIPC, PBGI, or LCB acquisition function yields cost-adjusted regret nearly matching the hindsight optimal. This shows our PBGI/LogEIPC stopping rule to be relatively robust against challenges in acquisition function optimization. Further discussions of our experiments across all cost types and cost-scaling factors can be found in Appendix˜D.
Figure 2: Cost-adjusted simple regret across acquisition-stopping rule pairs in 1D Bayesian regret setting. Objectives are sampled from a GP with Matérn-5/2 kernel, with a linear cost function scaled by $\lambda=0.1,0.01,0.001$. Matched PBGI pairs perform the best at $\lambda=0.1,0.01$, and remain close to hindsight-optimal at $\lambda=0.001$, though slightly outperformed by PRB and GSS paired with PBGI. The Immediate baseline is omitted at $\lambda=0.001$ due to its significantly higher regret (mean: 0.6942, error bar: [0.5314, 0.8570]) compared to other stopping rules.
Figure 3: Cost-adjusted simple regret across acquisition–stopping rule pairs in 8D Bayesian regret experiments. Objectives are sampled from a GP with a Matérn-5/2 kernel, using three different cost functions scaled by $\lambda=0.01$. Our PBGI/LogEIPC stopping rule consistently yields the lowest cost-adjusted regret that is close to hindsight optimal, when paired with LogEIPC, PBGI, and LCB.
@@ 4.2 Empirical Automated Machine Learning Benchmarks
We then evaluate on two empirical benchmarks based on use cases from hyperparameter optimization and neural architecture search: LCBench (Zimmer et al., 2021) and NATS-Bench (Dong et al., 2021).
LCBench provides training data of 2,000 hyperparameter configurations evaluated on multiple OpenML datasets. Due to the space constraints, we report results on the first three datasets from the lite version, covering image, mixed-type tabular, and large-scale numerical tabular classification tasks. In the unknown-cost experiments, the cost is the full (200-epoch) training time. For the known-cost setting, we observe that training time to scale approximately linearly with the number of model parameters. Based on a linear regression fit (see Appendix˜C), we adopt $10^{-3}\times p$, where $p$ is the number of model parameters, as a proxy of training time.
@@
We report the mean and error bars (2 times standard error) of the cost-adjusted simple regret, where we consider the evaluation costs to be some representative cost-scaling factor $\lambda$ (see Section˜3.2) multiplied with cumulative runtime. Simple regret is computed as the difference between (a) test error of the configuration with best validation error at the stopping time and (b) the best test error over all configurations. We present the results using proxy runtime here, and defer to Appendix˜D the additional results under (i) the cost model mismatch scenario (proxy runtime used during Bayesian optimization but actual runtime used for evaluation), and (ii) the unknown-cost scenario (actual runtime used during Bayesian optimization).
@@
Figure˜4 presents cost-adjusted regret for pairs of acquisition function and stopping rule on classification tasks from LCBench and NATS-Bench, with representative [^3] These were chosen to avoid degenerate cases—neither so large that the policy stops after only a few evaluations (e.g., $\lambda=10^{-4}$ on NATS-Bench, Figure 18) nor so small that evaluations are effectively free. cost-scaling factor $\lambda=10^{-4}$ and $\lambda=10^{-5}$ respectively. Additional results under different cost-scaling factors and the unknown-cost setting are provided in Figures˜15, 16 and 17 of Appendix˜D, along with visual comparisons between adaptive and fixed-iteration stopping. Overall, our PBGI/LogEIPC stopping rule consistently performs strongly in terms of cost-adjusted simple regret, particularly when paired with the PBGI acquisition function. We also report, in Table˜1 of Appendix˜D, how often each stopping rule fails to trigger within the 200-iteration cap: baselines such as SRGap-med and UCB–LCB often fail to stop early—particularly on NATS-Bench—whereas our stopping rule reliably stops before the cap.
@@
As we transition from model match to model mismatch, where the true objective function does not align perfectly with the Gaussian process prior, we find that our stopping rule, when paired with the PBGI acquisition function, continues to deliver performance close to the hindsight optimal, except on two NATS datasets likely affected by severe mismatch. In contrast, pairing with the LogEIPC acquisition function is less competitive. This degradation appears to stem from the relative advantage of PBGI over LogEIPC in higher dimensions or misspecified settings: as shown in Figure˜11 in Appendix˜D, PBGI maintains stronger performance under misspecification, thus contributing to the improved overall cost-adjusted regret. Thus, while our stopping rule is broadly robust, the choice of acquisition function remains an important consideration when facing objective model mismatch.
@@
Figure 4: Comparison of cost-adjusted simple regret across acquisition function and stopping rule pairs on LCBench and NATS-Bench. The objective is to minimize validation error on classification tasks, with proxy runtime as evaluation cost, scaled by representative values of $\lambda$ ($10^{-4}$ for LCBench and $10^{-5}$ for NATS-Bench); see Figures˜12, 13 and 14 for performance under other $\lambda$. The matched PBGI combination performs best on adult, higgs, and cifar10-valid, second-best on Fashion-MNIST, and third-best on the remaining two, while closely approaching the hindsight optimal across all tasks. Our stopping rule is also competitive when paired with LogEIPC or TS.

### v4

@@ 3.1 Theoretical Guarantees on Cost-Adjusted Simple Regret
Let $X$ be compact, and let $f:X\to\mathbb{R}$ be a random function with prior mean $\mu(\cdot)$. Consider a Bayesian optimization algorithm that begins at some initial point $x_{1}\in X$ with cost $C=c(x_{1})$, acquires subsequent points using either the PBGI or LogEIPC acquisition function, and terminates according to the PBGI/LogEIPC stopping rule. Let $\tau=\min_{t\geq 1}\{\sup_{x\in X}\alpha_{t}^{\mathrm{LogEIPC}}(x)<0\}$ be the algorithm’s stopping time, and denote the posterior expected improvement function by $\alpha_{t}^{\mathrm{EI}}(x)=\mathrm{EI}_{f\mid x_{1:t},y_{1:t}}\left(x;y^{*}_{1:t}\right)$. Then, for all $t<\tau$, $\alpha_{t}^{\mathrm{EI}}(x_{t+1})\geq c(x_{t+1})$.
As a consequence of this claim, we show in Theorem˜2 that the PBGI/LogEIPC stopping rule, when paired with PBGI or LogEIPC, achieves cost-adjusted simple regret no worse than stopping immediately after the initial evaluation. Although this may sound trivial, many acquisition function–stopping rule pairs fail to satisfy this no-worse-than-immediate property in practice (see Figures˜2 and 3), largely because most stopping rules are designed with simple regret alone in mind and do not explicitly account for evaluation costs. To our knowledge, this is the first formal guarantee on cost-adjusted simple regret for Bayesian optimization, generalizing beyond simpler settings such as Pandora’s Box, which assumes independent and discrete evaluations. All proofs in this section are in Appendix˜B.
@@
Consider the setting and algorithm specified in Lemma˜1. Let $U:=\mu(x_{1})-\mathbb{E}[\min_{x\in X}f(x)]<\infty$, then the algorithm’s expected cost-adjusted simple regret is bounded by
@@
This result yields a bound on expected cumulative cost and, under the natural assumption that practical evaluations have a positive minimum cost, a high-probability finite-time termination guarantee.
Corollary 3.
Consider the setting and algorithm specified in Lemma˜1. Then the expected cumulative cost of the algorithm is bounded by $\mathbb{E}\left[\sum_{t=1}^{\tau}c(x_{t})\right]\leq U+C$. Further, if evaluation costs are uniformly bounded below by a constant $c_{0}>0$, i.e., $c(x)\geq c_{0},\forall x\in X$, then for any $\delta\in(0,1)$, the algorithm terminates in at most $\frac{U+C}{\delta\cdot c_{0}}$ iterations with probability $1-\delta$.
@@ 3.1.1 Implication in the Budget-constrained Setting
We first note that in the discrete Pandora’s Box setting, under an expected budget constraint $B$, minimizing $\alpha_{t}^{\mathrm{PBGI}}(x)$ is Bayesian-optimal: it achieves the lowest expected simple regret, among all algorithms satisfying $\mathbb{E}\left[\sum_{t=1}^{\tau}c(x_{t})\right]\leq B$, and the cost function used in defining $\alpha_{t}^{\mathrm{PBGI}}(x)$ is $\lambda c(x)$ for some cost-scaling factor $\lambda$ which depends on $B$ (xie2024cost, Theorem 2).
@@
Consider the setting, algorithm, and notation specified in Lemmas˜1 and 2, but with costs rescaled by a factor $\lambda>0$: both the acquisition values and the stopping conditions are computed using $\lambda c(\cdot)$. If the cost-scaling factor is set to $\lambda=\frac{U}{B-C}$, then the algorithm’s expected cumulative unscaled cost satisfies $\mathbb{E}\!\left[\sum_{t=1}^{\tau}c(x_{t})\right]\leq B$.
If this choice of $\lambda$ proves overly conservative—leading to underspending within the budget—it can be paired with the PBGI-D variant of xie2024cost, which starts with $\lambda=\lambda_{0}$ and halves it each time stopping is triggered. Choosing $\lambda_{0}=U/(B-C)$ aligns the initial fixed-$\lambda$ phase with the budget in expectation, while ensuring that the adaptive decay is activated as designed. In Figure˜10 in Appendix˜D, we show PBGI-D with this choice of $\lambda_{0}$ is competitive.
@@ 3.2 Practical Implementation Considerations for PBGI/LogEIPC Stopping Rule
In practice, evaluation costs are often not known in advance. They can be modeled either (i) deterministically, using domain knowledge (e.g., proportional to model size in hyperparameter tuning), or (ii) stochastically, via a GP over log costs. In Section˜4, we present empirical benchmark results under both modeling approaches. For the stochastic approach, we follow astudillo2021multi to model $\ln c(x)$ as a GP with posterior mean $\mu_{\ln c}$ and variance $\sigma^{2}_{\ln c}$. To compute LogEIPC or PBGI, as well as their related stopping rules (including prior acquisition-based ones and ours), we replace $c(x)$ in Equations˜4, 5 and 9 with $\mathbb{E}[c(x)]=\exp(\mu_{\ln c}(x)+(\sigma_{\ln c}(x))^{2}/2)$. See Section˜D.3 for further discussion.
@@
Although our stopping rule has theoretical guarantees, in practice, it—like other adaptive stopping rules—can still suffer from spurious stops caused by two main sources. First, early in the optimization process, the GP model parameters often fluctuate significantly as new data points are collected, causing unstable stopping signals. To mitigate this, we enforce a stabilization period consisting of the first several evaluations, during which we do not allow any stopping rule to trigger. Second, imperfect optimization of the acquisition function—which is especially common in higher-dimensional search spaces—can lead to misleading stopping signals. To handle this, we use a debounce strategy, requiring the stopping rule to consistently indicate stopping over several consecutive iterations before stopping optimization. See Figure˜4 for an illustration.
@@ 4 Experiments
To evaluate our proposed PBGI/LogEIPC stopping rule, we design three complementary sets of experiments that progressively test its performance. First, we consider an idealized, low-dimensional Bayesian regret setting in which the GP model exactly matches the true objective. This controlled environment allows us to isolate the effect of different stopping rules without interference from modeling or optimization errors. Then, we move to a higher-dimensional Bayesian regret setting where each acquisition-function minimization must be approximated via a numerical optimizer. Finally, we test in practical scenarios with potential objective model mismatch, using the LCBench hyperparameter tuning benchmark and the NATS neural architecture size search benchmark. In each case, we compare pairs of acquisition functions and stopping rules, which we describe next.
@@
We consider four common acquisition functions that were discussed in Section˜2: LogEIPC, PBGI, LCB, and TS—chosen for their competitive performance, computational efficiency, and close connections to existing stopping rules.
@@
We compare the proposed PBGI/LogEIPC stopping rule against several stopping rules from prior work: UCB–LCB (makarova2022automatic), LogEIPC-med (ishibashi2023stopping), SRGap-med (ishibashi2023stopping), and PRB (wilson2024stopping). We also include two simple heuristics used in practice: Convergence and GSS. UCB–LCB stops once the gap between upper and lower confidence bounds falls below a configurable threshold $\theta$. LogEIPC-med stops when the log expected improvement per cost drops beneath $\log(\eta)$ plus the median of its initial $I$ values. SRGap-med stops when the gap of the expected minimum simple regret falls below $\chi$ times the median of its initial $I$ values. PRB triggers once a probabilistic regret bound satisfies the regret tolerance $\epsilon$ and confidence $\delta$ parameters. Convergence stops as soon as the best observed value remains unchanged for $w$ iterations. GSS stops if the recent improvement is less than $\phi\times\mathrm{IQR}$ over the past $w$ trials where $\mathrm{IQR}$ denotes the inter-quartile range of current observations. Finally, we include two reference baselines: Immediate, which stops right after the initial evaluation (see Section˜3.1), and Hindsight, which, for each trial, selects the stopping time that yields the lowest cost-adjusted simple regret in hindsight, thus providing a lower bound on achievable performance.
@@ 4.1 Bayesian Regret
In the 1-dimensional experiments, we perform an exhaustive grid search over $[0,1]$ to isolate the effect of stopping behavior from numerical optimization. In the 8-dimensional setting, due to instability from higher dimensionality, we apply moving average with a window size of 20 when computing the PBGI/LogEIPC stopping condition. See Figure˜4 in Appendix˜C for an illustration. More details on our experiment setup, and computational considerations are provided in Appendices˜C and D.
Figure˜2 shows a comparison of cost-adjusted regret for acquisition function and stopping rule pairings under different cost-scaling factors $\lambda$ in the 1-dimensional setting and under different cost regimes (uniform, linear, periodic) in 8-dimensional setting.
In the 1-dimensional setting, the Bayesian optimization problem is relatively straightforward and all acquisition functions attain nearly identical hindsight optima, independent of cost scale or cost type. From the plot, our PBGI/LogEIPC stopping rule pairing with LogEIPC or PBGI acquisition function delivers competitive performance for each $\lambda$, and is particularly strong in handling high-cost scenarios—those with large $\lambda$.
In the 8-dimensional experiments, however, clear gaps emerge: Under uniform and linear costs, LogEIPC, PBGI, and LCB perform similarly and substantially outperform TS, while in the periodic case, LogEIPC and PBGI outperform LCB and TS. In every cost regime, combining our stopping rule with the LogEIPC, PBGI, or LCB acquisition function yields cost-adjusted regret nearly matching the hindsight optimal. This shows our PBGI/LogEIPC stopping rule to be relatively robust against challenges in acquisition function optimization. Further discussions of our experiments across all cost types and cost-scaling factors can be found in Appendix˜D.
Figure 2: Cost-adjusted simple regret across acquisition-stopping rule pairs in 1D and 8D Bayesian regret setting. In 1D, objective functions are sampled from a GP with a Matérn-5/2 kernel and a linear cost function scaled by $\lambda=0.1,0.01,0.001$. The Immediate baseline is omitted at $\lambda=0.001$ due to its much higher regret (mean 0.6942, error bar [0.5314, 0.8570]). In 8D, objective functions are also drawn from a GP with a Matérn-5/2 kernel, using three cost functions scaled by $\lambda=0.01$.
@@ 4.2 Empirical Automated Machine Learning Benchmarks
We then evaluate on two empirical benchmarks based on use cases from hyperparameter optimization and neural architecture search: LCBench (ZimLin2021a) and NATS-Bench (dong2021nats).
LCBench provides training data of 2,000 hyperparameter configurations evaluated on 35 OpenML vanschoren2014openml datasets. In the unknown-cost experiments, the cost is the full (200-epoch) training time. For the known-cost setting, we observe that training time to scale approximately linearly with the number of model parameters. Based on a linear regression fit (see Appendix˜C), we adopt $\alpha p$ as a proxy of training time, where $p$ is the number of model parameters and $\alpha$ is a dataset-specific coefficient estimated from the regression.
@@
We report the mean and error bars (2 times standard error) of the cost-adjusted simple regret, where we consider the evaluation costs to be some representative cost-scaling factor $\lambda$ (see Section˜3.2) multiplied with cumulative runtime. Simple regret is computed as the difference between (a) test error of the configuration with best validation error at the stopping time and (b) the best test error over all configurations. We present the results using proxy runtime here, and defer to Appendix˜D the additional results under (i) the cost model mismatch scenario (proxy runtime used during Bayesian optimization but actual runtime used for final performance evaluation), and (ii) the unknown-cost scenario (actual runtime used during Bayesian optimization).
@@
Figure˜3 presents performance of acquisition function–stopping rule pairs on LCBench and NATS-Bench. For LCBench, we report min–max normalized cost-adjusted simple regret (see definition in Appendix˜D) aggregated across 35 datasets in Figure˜3 evaluated under three representative [^4] These $\lambda$ values are chosen to avoid degenerate cases—neither so large that the policy stops after only a few evaluations (e.g., $\lambda=10^{-4}$ on NATS-Bench, see Figure 17) nor so small that evaluations are effectively free. cost-scaling values. For NATS-Bench, we report cost-adjusted simple regret under $\lambda=10^{-5}$, with additional results under two other cost-scaling factors shown in Figures˜14, 15 and 16 in Appendix˜D.
Per-dataset LCBench results are provided in Figures˜19, 20 and 21 in the appendix. Across these 35 tasks, around 75% of them show our PBGI/LogEIPC stopping rule performing competitively when paired with either PBGI or LogEIPC. The remaining outliers are mostly (aside from two exceptions) very small datasets with <10000 instances that might lead to severe model misspecification.
Additional results under the cost model mismatch setting and the unknown-cost setting are provided in Figures˜17 and 18 in Appendix˜D, along with comparisons between our adaptive stopping rule and fixed-iteration baselines. Overall, our PBGI/LogEIPC stopping rule consistently performs strongly in terms of cost-adjusted simple regret, when paired with either PBGI or LogEIPC. We also report, in Table˜1 of Appendix˜D, how often each stopping rule fails to trigger within the 200-iteration cap: baselines such as SRGap-med and UCB–LCB often fail to stop early—particularly on NATS-Bench—whereas our stopping rule reliably stops before the cap.
@@
As we transition from model match to model mismatch, where the true objective function does not align perfectly with the GP prior, we find that our stopping rule, when paired with the PBGI acquisition function, continues to deliver performance close to the hindsight optimal, except on two NATS datasets likely affected by severe mismatch. In contrast, pairing with the LogEIPC acquisition function is less competitive. This degradation appears to stem from the relative advantage of PBGI over LogEIPC in higher dimensions or misspecified settings: as shown in Figure˜10 in Appendix˜D, PBGI maintains stronger performance under misspecification, thus contributing to the improved overall cost-adjusted regret. Thus, while our stopping rule is broadly robust, the choice of acquisition function remains an important consideration when facing objective model mismatch.
@@
Figure 3: Cost-adjusted simple regret across acquisition function–stopping rule pairs on LCBench and NATS-Bench. The objective is to minimize validation error on classification tasks, with scaled proxy runtime as evaluation cost, scaled by representative values of $\lambda$ ($10^{-3},10^{-4},10^{-5}$ for LCBench and $10^{-5}$ for NATS-Bench). For LCBench, results are aggregated across 35 datasets using min–max normalization. Our PBGI/LogEIPC stopping rule, when paired with either LogEIPC or PBGI, typically ranks among the top 3 pairs and closely approaches the hindsight optimal on LCBench and on cifar10-valid in NATS-Bench, but slightly worse on the other two NATS datasets.